Q2.Calculate the current I by superposition. 5A 122 2or 40v 252 you 7.5 A tv You
Answer #1

the Circuit as Solution) We are given Shown and we are required find the worent. I to 5A 122 I M 20.2 4or w 7 40V + 252 7.5A2001 122 applying sourne parallel weget this looksourue in 100 V 2 100 V L I 2 5A Now AAN w source 402 252 transformation und20.2 wir Iz sor thus I=/ R Eur 402 225-32 +) 300 V R- 40+20+ 251112 402 THO =100+ 8.108 108.108 = 2.775 A 5 I = 300 108.108 (​​​​​​​Therefore required value was calculated out using superposition thereom and each step was well explained